CHW 1: Ch 20 1-3 )5 of 7 where K = eo = 8.854 × 10-12 C2/N ㎡ ) is the permittivity of free space Consider two point charges located on the x axis: #1 -1.750 m ; the second charge, 92.38.0 nC =-16.5 nC , is located at is at the origin(-0.0000)
Answer #1

d 2 .12 4x3.1#x3.g 54 x 16-12 dl 2 χご -15m -1.225mm 13 1-38hc 55nc = 55x10-3 c Co.53)d2 = O + 1,225-1.225 n .225 2 2.3 13 23 The Net Force (-2.91-1.25) x10-S て Negative sign shown the dine tloA。f Shouh 1 S